Project

General

Profile

NDOS Commissioning Meeting 2011/09/08

Attendance: Jon, Mark Messier, Martin, Susan, Nick, Jarek, Jan, Minerba, Peter, Mark Mathis, Chad, Zukai, Athans, Gavin,

Agenda:

  1. Beam & Detector Update
  2. DataCheck Update
  3. DAQ Update
  4. DCS Update
  5. BNB Event Update
  6. DSO on multiple channels simultaneously
  7. AOB

Beam & Detector Update

  • Jon: pretty stable this past week.
  • Jon: Chuck Andrews sent a short laundry list of minor improvements that could be made to the grounding out at NDSB. Jon will follow up with Rick T.
  • Mark: status of new target? Jon will investigate.
  • Chad: BNB timing peak is still present and stable.
  • Mat: 6 APDs have died over the last week. Will try to run these warm.

DataCheck Update

  • NuMI peak looks good.
  • Some empty events in some runs, but likely towards the end of the run or at the beginning of the run.
  • nearline:
    • processing new subruns every ~90 minutes.
    • webpage is updated every 5 minutes.
    • NuMI timing peak is now part of the standard plots
    • cosmic muon rate is also now part of the standard plots
    • Jon will work with Jarek on putting a shift checklist together.
    • DDS is running on farm-node-16 for some reason, and the memory leak brought the machine down twice last week. Andrew N. & co. will investigate.

DAQ Update:

  • four DAQ software packages to be updated tomorrow morning: ConfigurationManager, DatabaseUtils, EventDispatcherCommandSet, GlobalTrigger

DCS Update

  • Martin:
    • a timestamp has been added to the overview so that one can tell that the monitor is still running
    • changed some alarm thresholds for an APD so now everything is "green", so if anything goes red, it actually does mean there is a problem

AOB

  • Chris Backhouse is ~ready to run his special DSO code on the detector. Details of how this is to be done will be worked out between him and the DAQ software group. Jon will initiate the conversation.